If 3 red-yellow tokens are flipped, what is the theoretical probability of all 3 tokens landing with the same color up?
Answer by flame8855(424)

all red : 1/2*1/2*1/2= 1/8
all yellow : 1/2*1/2*1/2= 1/8
total= 1/8 + 1/8 = 2/8 =1/4